Circular No. 7334

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999gm IN PGC 24106 M. Papenkova and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova in PGC 24106. SN 1999gm was discovered and confirmed on unfiltered images taken on Dec. 15.4 and 16.4 UT, both showing mag ∼ 17.3. The new object is located at α = 8^{h}35^{m}11^{s}.73, δ = −8^{o}23′15′′.4 (equinox 2000.0), which is 12′′.1 east and 1′′.5 south of the nucleus of PGC 24106. A KAIT image of the same field taken on Dec. 5.4 already showed a hint of the supernova. A KAIT image taken on Nov. 24.4 (limiting mag ∼ 19.0) showed nothing at the position of SN 1999gm.

D. C. Leonard, R. Chornock, A. L. Coil, and A. V. Filippenko, University of California at Berkeley, report that preliminary reduction of CCD spectra (range 330–1000 nm), obtained on Dec. 17 UT with the Shane 3-m reflector at Lick Observatory, shows that SN 1999gm is a type-Ia supernova near maximum light.

SUPERNOVA 1999gi IN NGC 3184 Unfiltered CCD magnitudes: Dec. 3.18 UT, [18.5 (O. Trondal, Oslo,

Norway; communicated by B. H. Granslo); 10.73, 14.3 (R. Kushida, Yatsugatake South Base Observatory; communicated by S. Nakano); 11.794, 14.6 (S. Yoshida and K. Kadota, Ageo, Saitama, Japan).

COMET 141P/MACHHOLZ 2 Total m_{1} estimates of components A and D, respectively (visual unless

otherwise noted): Nov. 12.70 UT, 13.3, 12.0 (M. Jäger, near Vienna, Austria, 0.25-m Schmidt camera + Kodalith orthochromatic film); 26.68, 11.8, 12.3 (Jäger); 27.22, 11.6, 12.3 (N. Biver, Oahu, HI, 0.26-m reflector); 27.73, 12.8, 12.9 (W. Hasubick, Buchloe, Germany, 0.44-m reflector); 28.22, 12.1, 12.6 (Biver); 29.71, 11.8, 12.0: (M. Meyer, Frauenstein, Germany, 0.25-m reflector); 30.40, 13.1, 14.4: (A. Nakamura, Kuma, Ehime, Japan, 0.60-m reflector + CCD); Dec. 2.45, 11.4, 12: (D. A. J. Seargent, Cowra, N.S.W., 0.25-m reflector); 3.40, 12.5, 13.6: (Nakamura); 4.70, 10.6, 11.8: (Meyer); 5.48, 12.0:, [12.0 (M. Mattiazzo, Wallaroo, S. Australia, 0.20-m reflector); 6.79, 11.8, 13.0 (J. Carvajal, Avila, Spain, 0.45-m reflector); 8.40, 11.9, 14.0: (Nakamura).

1999 December 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green